OpenStudy (anonymous):

Find dy/dx, using implicit differentiation. 4x + 7y = xy dy dx= Compare your answer with the result obtained by first solving for y as a function of x and then taking the derivative. y= dy dx= I am stuck on this problem

OpenStudy (mani_jha):

If you take dy/dx, you get; 4(dx/dx)+7(dy/dx)=d(xy)/dx 4+7(dy/dx)=x(dy/dx)+y(dx/dx) 4+7(dy/dx)=x(dy/dx)+y dy/dx(7-x)=y-4 dy/dx=(y-4)/(7-x) Just find in the original equation, y in terms of x. Now, can you do it?

OpenStudy (anonymous):

one second

OpenStudy (anonymous):

4(dx/dx) that first dx is dy correct?

OpenStudy (mani_jha):

No, we are finding the derivate of the whole thing with respect to x. So, whatever there is in the numerator, we will differentiate it with respect to x. (dx/dx) will be 1

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

okay so fir this part x(dy/dx)+y(dx/dx) why is the y(dx/dx) not y(dy/dx)?

OpenStudy (anonymous):

4x + 7y = xy y': 4 + 7y' = y + xy' x': 4x' + 7 = x'y + x

OpenStudy (anonymous):

In plain version :)

OpenStudy (anonymous):

y': 4 + 7y' = y + xy' that would be the only one i need correct instead of the x'?

OpenStudy (anonymous):

Compare your answer with the result obtained by first solving for y as a function of x and then taking the derivative. y =

OpenStudy (anonymous):

how would i do that step now?

OpenStudy (anonymous):

What did you come up?

OpenStudy (anonymous):

y = 4 + 7y/x

OpenStudy (anonymous):

Solving for y as function of x

OpenStudy (anonymous):

4x/x-7 is what i got and it was correct

OpenStudy (anonymous):

but i have no idea how to get that answer...

OpenStudy (anonymous):

For the second part?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

Let me do on paper first!

OpenStudy (anonymous):

dy/dx=(y-4)/(7-x) this was the answer to the first part

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

I don't see where to get the denominator ( 7-x)?

OpenStudy (anonymous):

y': 4 + 7y' = y + xy' 7y' - xy' = y -4 ->( 7 -x) y' = y -4 => y' = ( y - 4)/ ( 7 -x)

OpenStudy (anonymous):

4x+7y=xy is the original equation so you move the xy to the side with 7y xy-7y=4x then you factor out a y y(x-7)=4x then you divide by x-7 y=4x/(x-7)

OpenStudy (anonymous):

I forgot we have to factor it out y' first :)

OpenStudy (anonymous):

If you take derivative of y = 4x/(x-7) I can't find the match with the result of first answer :(

OpenStudy (anonymous):

@Mani Jha, give some clue for the second part will you!

OpenStudy (anonymous):

y = 4 + 7y/x ??

OpenStudy (mani_jha):

The second part should give the same answer as the first. Wait a second

OpenStudy (anonymous):

Oh, I got it

OpenStudy (anonymous):

I guess because I'm sort of relyng on you

OpenStudy (anonymous):

y = 4x/ ( 7 -x) y' = 4 (7-x) - 4x/ ( 7-x)^2

OpenStudy (mani_jha):

That should be x-7 downstairs, isnt it?

OpenStudy (anonymous):

= 4 ( 7-x)/ (7-x)^2

OpenStudy (anonymous):

Cancel out

OpenStudy (mani_jha):

y=4x/(x-7) y'={(x-7)4-4x}/(x-7)^2 =-28/(x-7)^2 Now we substitute y=4x/(x-7) in our first result we get -28/(x-7)^2, which is same I still dont see how you get 7-x downstairs!

OpenStudy (anonymous):

(28 - 8x)/ (7-x)^2

OpenStudy (mani_jha):

you substituted y=4x/(x-7), right? y' = {(4x)/(x-7)-4}/(7-x) ={(4x-4x+28)/(7-x)(x-7)} =28/(7-x)(x-7) =-28/(x-7)^2

OpenStudy (anonymous):

OH, that's my problem. All I do is take derivative!

OpenStudy (anonymous):

@camaron, do you understand what Mani has showed? @Mani, thanks a lot! I learn the new substitution today :)

OpenStudy (mani_jha):

No problem, dude ;)